# GATlab: Modeling and Programming with Generalized Algebraic Theories

### v1

@@ 1.2 Related work
GATlab’s emphasis on theory morphisms as a means of translating between theories and combining theories via colimits is also found in the mathematical knowledge management system MMT [32], which supports more general classes of theories but is less focused on computational semantics.
@@ 2 Background: Generalized algebraic theories and their models
We review the main ideas behind generalized algebraic theories (GATs), focusing on their syntax, in both mathematical and GATlab programming notation, and their standard set-theoretic semantics. GATs were introduced by John Cartmell in his PhD thesis [6] and later in print [7], to which we refer for the many details omitted here. As further references, Pitts [31, §6] and Taylor [36, Chapter VIII] present type theories based closely on Cartmell’s GATs.
@@ 2.1 Syntax of GATs
GATs are most easily understood through examples. Figure 1 shows the ur-example, the theory of categories, first in mathematical notion in the style of Pitts [31] and then as a theory in the GATlab programming environment. Most notably, the listings are the same up to cosmetic differences in syntax. Julia’s macro system makes this possible by capturing the code in the body of the @theory macro call instead of evaluating it as standard Julia code. Also notice that GATlab allows text identifiers to be aliased by mathematical symbols (via Unicode operators) and axioms to be tagged with names.
@@ 2.2 Semantics of GATs
Under the standard set-theoretic semantics, a model of a GAT interprets each type constructor with a set, or family of sets, or family of families of sets, etc., according to the level of type dependency; and similarly each term constructor with a function, or family of functions, etc., according to the level of dependency; in such a way that all of the term equalities are interpreted as true equations. A more precise account is given by Cartmell [7, §11], but the idea becomes most transparent in examples. A model $M$ of the theory of categories (Figure 1) consists of a set of objects, $M(\operatorname{Ob})$; a family of hom-sets, $(M(\operatorname{Hom})(a,b))_{a,b\in M(\operatorname{Ob})}$; the identities, a family of functions $(M(\operatorname{id})(a):1\to\operatorname{Hom}(a,a))_{a\in M(\operatorname{Ob})}$; and finally the composition operations, a family of functions
@@
GATlab offers the programmer two different ways to work with models of GATs, as we will see. Generalizing their set-theoretic semantics, models of GATs can be taken in any contextual category [7, §14]. We do not elaborate on the categorical semantics of GATs as it plays no role in GATlab.
@@ 2.3 Alternatives to GATs
Generalized algebraic theories belong to a family of essentially equivalent logics that extend the logic of algebraic theories to encompass the theory of categories and others like it. Besides GATs, the best known of these logics are essentially algebraic theories [1, §3.D], finite limit sketches, and finite limit theories. Cartmell sketched an argument to the effect that, under their set-theoretic semantics, generalized and essentially algebraic theories have equivalent categories of models [7, §6]. Meanwhile, essentially algebraic theories and finite limit sketches are directly intended to be syntactic presentations of finite limit theories, an invariant notion of theory.
@@ 3.1 Models of algebraic theories
In order to implement module-style models using Julia’s multiple dispatch, we employ a trick. We associate model implementations with structs (records in Julia), and then use those structs to guide dispatch. Following the module-style, the model of the theory of monoids with integer addition takes the following form.
@@
We also a provide a convenience macro @withmodel for evaluating expressions in a module-style module. For example, the invocation
@@ 3.2 Incorporating dependent types into models
We can recover the fibered approach from the indexed one by letting $\operatorname{Hom}^{\prime}=\operatorname{IsHom}^{-1}(\mathbf{true})$, but there is no systematic way to “drop” redundant information, so in a sense the second approach is more primitive than the first. This is related to the fact that the Grothendieck construction only gives an equivalence of categories, not an isomorphism. [^7] One might also think about the indexed approach as a form of the “type erasure” that happens during the compilation of dependently typed languages, where values not relevant to the runtime semantics are erased. Perhaps a sufficiently intelligent compiler could in fact erase the redundant copies of objects in a diagram, but the Julia compiler does not work like that.
@@ 3.3 Free models of GATs
This code creates a new module called FreeCategory in which structs called Ob and Hom are defined and methods ThCategory.id and ThCategory.compose implemented for them. We can use the resulting data structures as follows.
@@ 4.1 Syntax of GAT morphisms
Following Cartmell [7, §12], two GATs $A$ and $B$ can be related to each other via an interpretation, which assigns to each type (respectively, term) constructor of $A$ a type (respectively, term) in $B$. This mapping must satisfy some conditions in order for it to be a well-typed and validity-preserving translation of the vocabulary of $A$ into the vocabulary of $B$.
@@ 4.2 Data structures for subclasses of GAT morphisms
Morphisms of GATs are highly expressive, which entails a certain amount of complication. It can be useful to consider GAT morphisms that are more restrictive, yet require less data to specify and have better computational affordances. GATlab defines no less than four data structures that implement GAT morphisms of varying expressivity:
@@ 4.3 Computation with GAT morphisms
Pushing forward types and terms along SimpleTheoryMap or TheoryMap instances requires an implementation of pattern matching and syntactic substitution. The pushforward operation is essential to pulling back models along theory maps: models associate a computational semantics with each expression of a theory $B$, so by precomposing with the interpretation function of expressions of $A$ to $B$, we see that GAT morphisms $A\rightarrow B$ induce a function from models of $B$ to models of $A$. For example, if ThArith were given a model JuliaArith, using the native addition and multiplication of Julia, then the model IntPlusMonoid defined manually in the previous section could be obtained declaratively by calling GATlab’s migrate_model on PlusM and JuliaArith. As another example, the following theory map allows us to automatically generate models of opposite categories in a manner that is declarative and statically verifiable:
@@
The operation of checking the validity of a theory map is trivial for IdTheoryMap. For InclTheoryMap, one must check that the identifiers found in dom are a strict subset of those found in codom. For SimpleTheoryMap, one checks that for each type and term constructor of the domain that the assigned type or term in the codomain has the same structure. E.g., to check the assignment trans => compose, one looks at the generic term for trans, namely
@@
compose(p,q) $\dashv$ [(a,b,c)::Ob, p::(a $\to$ b),q::(b $\to$ c)]
For a general theory map, determining validity-preservation requires solving many word problems, and deciding equality is a nontrivial semi-decidable computation in general. For example, one must check that SwapMonoid preserves the associativity of monoid multiplication.
@@
5 Implementation of GATlab
@@ 5.1 Scopes
While this seems like a relatively minor detail, it was not trivial to settle on this approach to scoping, and so we hope that reporting the details of this scope solution may help language designers in the future.
5.2 Expanding macros inside macros
@@
6 Applications and Extensions
  1. (1)
  2. (2)
    fibered vs. indexed point of view on dependent types (section 3.2)
  3. (3)
    hygenic substitution via scope tags (section 5.1)
  • Systems can be optimized/compiled to produce more efficient code
  • Systems can be transferred between different programming languages
{ack}

### v2

@@ 1.2 Related work
GATlab’s emphasis on theory morphisms as a means of translating between theories and combining theories via colimits is also found in the mathematical knowledge management (MKM) system MMT [34], which supports more general classes of theories but is less focused on computational semantics. Likewise, this approach to MKM is found in MathScheme [6] and related work on theory presentation combinators [7].
@@ 2 Background: Generalized algebraic theories and their models
We review the main ideas behind generalized algebraic theories (GATs), focusing on their syntax, in both mathematical and GATlab programming notation, and their standard set-theoretic semantics. GATs were introduced by John Cartmell in his PhD thesis [8] and later in print [9], to which we refer for the many details omitted here. As further references, Pitts [33, §6] and Taylor [38, Chapter VIII] present type theories based closely on Cartmell’s GATs.
@@ 2.1 Syntax of GATs
GATs are most easily understood through examples. Figure 1 shows the ur-example, the theory of categories, first in mathematical notion in the style of Pitts [33] and then as a theory in the GATlab programming environment. Most notably, the listings are the same up to cosmetic differences in syntax. Julia’s macro system makes this possible by capturing the code in the body of the @theory macro call instead of evaluating it as standard Julia code. Additionally, GATlab allows text identifiers to be aliased by mathematical symbols (via Unicode operators) and axioms to be tagged with names. Note that we break with convention in type theory and put the context on the right so that the name of the constructor is the first thing that one sees on a given line.
@@ 2.2 Semantics of GATs
Under the standard set-theoretic semantics, a model of a GAT interprets each type constructor with a set, or family of sets, or family of families of sets, etc., according to the level of type dependency; and similarly each term constructor with a function, or family of functions, etc., according to the level of dependency; in such a way that all of the term equalities are interpreted as true equations. A more precise account is given by Cartmell [9, §11], but the idea becomes most transparent in examples. A model $M$ of the theory of categories (Figure 1) consists of a set of objects, $M(\operatorname{Ob})$; a family of hom-sets, $(M(\operatorname{Hom})(a,b))_{a,b\in M(\operatorname{Ob})}$; the identities, a family of functions $(M(\operatorname{id})(a):1\to\operatorname{Hom}(a,a))_{a\in M(\operatorname{Ob})}$; and finally the composition operations, a family of functions
@@
GATlab offers the programmer two different ways to work with models of GATs, as we will see. Generalizing their set-theoretic semantics, models of GATs can be taken in any contextual category [9, §14]. We do not elaborate on the categorical semantics of GATs as it plays no role in GATlab.
@@ 2.3 Alternatives to GATs
Generalized algebraic theories belong to a family of essentially equivalent logics that extend the logic of algebraic theories to encompass the theory of categories and others like it. Besides GATs, the best known of these logics are essentially algebraic theories [1, §3.D], finite limit sketches, and finite limit theories. Cartmell sketched an argument to the effect that, under their set-theoretic semantics, generalized and essentially algebraic theories have equivalent categories of models [9, §6]. Meanwhile, essentially algebraic theories and finite limit sketches are directly intended to be syntactic presentations of finite limit theories, an invariant notion of theory.
@@ 3.1 Models of algebraic theories
In order to implement module-style models using Julia’s multiple dispatch, we employ a trick. We associate model implementations with structs (records in Julia), and then use those structs to guide dispatch. In the following code, the Tuple{Int} parameter to Model tells the system that when IntPlusMonoid is used as a model of a theory, the first type constructor in the theory should be interpreted as Int. We will discuss why we do this later on; in brief this “exposes” the implementing types to the Julia type system, which makes it easier to create derived models.
@@
We also a provide a convenience macro @withmodel for evaluating expressions in a module-style model. For example, the invocation
@@ 3.2 Incorporating dependent types into models
We can recover the fibered approach from the indexed one by letting $\operatorname{Hom}^{\prime}=\operatorname{IsHom}^{-1}(\mathbf{true})$, but there is no systematic way to “drop” redundant information, so in a sense the second approach is more primitive than the first. This is related to the fact that the Grothendieck construction only gives an equivalence of categories, not an isomorphism. [^7] One might also think about the indexed approach as a form of the “type erasure” that happens during the compilation of dependently typed languages, where values not relevant to the runtime semantics are erased. That is, the runtime representation of values of type Hom(a,b) need not carry around pointers to a and b. Ideally, this would be an optimization performed by a clever compiler rather than done manually, but we must work within the constraints of the Julia programming language.
@@ 3.3 Free models of GATs
This code creates a new module called FreeCategory in which structs called Ob and Hom are defined and methods ThCategory.id and ThCategory.compose implemented for them. It also provides some special logic via an externally-defined function associate_unit to enforce that composition is strictly associative and unital, by normalizing to $n$-ary composition operations whenever compose is called. We can use the resulting data structures as follows.
@@
Currently, we have implemented the following normalization strategies:
  • associate, normalizing a binary operation for associativity
  • associate_unit, normalizing a binary operation for associativity and unitality
  • associate_unit_inv, normalizing a binary operation for associativity, unitality, and inverses
  • distribute_unary, distribute a unary operation over a binary operation
  • involute, normalize an involutive unary operation
  • normalize_zero, collapse to zero when the expression contains a zero
However, these are simply the normalization strategies that we have found useful; our approach does not preclude other strategies from being implemented.
@@ 4.1 Syntax of GAT morphisms
Following Cartmell [9, §12], two GATs $A$ and $B$ can be related to each other via an interpretation, which assigns to each type (respectively, term) constructor of $A$ a type (respectively, term) in $B$. This mapping must satisfy some conditions in order for it to be a well-typed and validity-preserving translation of the vocabulary of $A$ into the vocabulary of $B$.
@@ 4.2 Data structures for subclasses of GAT morphisms
Morphisms of GATs are highly expressive, which entails a certain amount of complication. It can be useful to consider GAT morphisms that are more restrictive, yet require less data to specify. These more restricted GAT morphisms also make it easier to compute, for instance, pushouts. GATlab defines no less than four data structures that implement GAT morphisms of varying expressivity:
@@ 4.3 Computation with GAT morphisms
Pushing forward types and terms along SimpleTheoryMap or TheoryMap instances requires an implementation of pattern matching and syntactic substitution. The pushforward operation is essential to pulling back models along theory maps: models associate a computational semantics with each expression of a theory $B$, so by precomposing with the interpretation function of expressions of $A$ to $B$, we see that GAT morphisms $A\rightarrow B$ induce a function from models of $B$ to models of $A$. For example, if ThArith were given a model JuliaArith, using the native addition and multiplication of Julia, then the model IntPlusMonoid defined manually in the previous section could be obtained declaratively by calling GATlab’s migrate_model on PlusM and JuliaArith. As another example, the following theory map allows us to automatically generate models of opposite categories. Once we have checked that the theory map is valid, we have confidence that applying it to correct models will always produce correct models.
@@
The operation of checking the validity of a theory map is trivial for IdTheoryMap. For InclTheoryMap, one must check that the identifiers found in dom are a strict subset of those found in codom. For SimpleTheoryMap, one must check that for each type and term constructor of the domain that the assigned type or term in the codomain has the same structure. E.g., to check the map F’s assignment compose => tran, one looks at the domain’s generic term in context, namely
compose(f,g) $\dashv$ [(a,b,c)::Ob, f::(a $\to$ b), g::(b $\to$ c)]
@@
For a general theory map, determining validity-preservation requires solving many word problems, and deciding equality is a nontrivial semi-decidable computation in general. For example, one must check that OpMonoid preserves the associativity of monoid multiplication.
@@
5 Applications and Extensions
  1. (1)
  2. (2)
    fibered vs. indexed point of view on dependent types (section 3.2)
  3. (3)
    hygenic substitution via scope tags (section A.1)
  • Systems can be optimized/compiled to produce more efficient code
  • Systems can be transferred between different programming languages
{ack}
Appendix A Implementation of GATlab
@@ A.1 Scopes
While implementing lexical scopes might seem a minor detail, it was not trivial to settle on this approach to scoping, and so we hope that reporting the details of this solution will help designers of typed computer algebra systems in the future. Using a combination of UUID-based scope identifiers and consecutive-integer indexing is more ergonomic than pure deBruijn indexing because reindexing happens less frequently. More importantly, it is also better at detecting programmer errors. In a computer algebra system, we are frequently moving syntax trees around, so it may happen that we forget to bring the corresponding scope along; having variables reference their scope by UUIDs is an easy way to detect this kind of error.
A.2 Expanding macros inside macros